OpenStudy (anonymous):

Vector question: Given: (aXb)*c=1 Where a, b and c are vectors, X is vector multiplication and * is scalar multiplication. u=a+b v=b+c w=a+c Express (uXv)*w in terms of a. b and c in the simplest possible way.

OpenStudy (anonymous):

@experimentX @dumbcow @ash2326 @apoorvk

OpenStudy (anonymous):

((a+b)X(b+c))*(a+c)=?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

look at the integer next to it "1"

OpenStudy (experimentx):

just simplify it. ((a+b)X(b+c))*(a+c) (axb + axc +bxb + bxc )*(a+c) (axb + axc + 0 + bxc )*(a+c) (axb*a + axc*a + bxc*a + axb*b + axc*b + bxc*b ) (0 + 0 + bxc*a + 0 + axc*b + 0 ) a*bxc + a*cxb = a*bxc + (-a*bxc) = 0

OpenStudy (anonymous):

In the end that's pretty much what I did, but I ended up with a different result. I just wasn't sure about how to multiply out cross products. Thanks